OpenStudy (anonymous):

conservation of momentum... how do i find final velocity given initial mass 1, initial mass 2, and initial velocities of 1 and 2 ?

OpenStudy (anonymous):

i tried using the equations m1v1 + m2v2 = m1v1' + m2v2' and 1/2 m1v1^2 etc (conservation of kinetic energy). i solved for v1' and plugged it into the conservation equation to solve for v2' and the algebra was just super hairy... am i even doing this right?

OpenStudy (shane_b):

It would be easier to explain if you posted the actual problem and if possible, show what you've tried :)

OpenStudy (anonymous):

two carts with masses m1=.8kg and m2=1.2 kg are moving toward each other with speeds v1=4.5m/s and v2= 3.8 m/s. what is the final speed of m1?

OpenStudy (anonymous):

i got v1'= (m1v1 + m2v2 - m2v2')/m1 from the conservation of momentum equation and v2' = sqrt of (m1v1^2 + m2v2^2 - m1v1'^2)/m2 from the conservation of kinetic energy

OpenStudy (anonymous):

if the collision is elastic then u1 - u2 = v2-v1

OpenStudy (anonymous):

\[V1 = \frac{ U1(M1-M2) +2M2U2 }{ m1+m2 }\]

OpenStudy (anonymous):

is v1 final velocity?

OpenStudy (anonymous):

Yup...